Please help me put this into an equation. A man has $.70 in dimes and nickels. He has 8 coins together. How many nickels does he have? I know the answer is 2 by trial and error but I don't know the equation.
Found 2 solutions by Paul, checkley71:
Answer by Paul(988)

Let the dimes be x
Let the nickels be y
x+y=8
y=8-x (subsitution)
DDime value x 100 = 10
Nickel value x 100 = 5
TOtal value x 100 = 70

Equation:
10x+5y=70
10x+5(8-x)=70
10x-5x=70-40
5x=30
x=6

y=8-6
y=2

Hence, there are 6 dimes and 2 nickels.
Paul.

Answer by checkley71(8403)

5X+10(8-X)=70 OR 5X+80-10X=70 OR -5X=110 OR 5X=10 OR X=10/5 OR X=2 NICKELS
PROOF 5(2)+10(8-2)=70 OR 10+10(6)=70 OR 10+60=70